OpenStudy (anonymous):

Show all work to solve 2/x + 1/9 =1/3

OpenStudy (anonymous):

do you know taking LCM

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

2/x = 1/3 -1/9 2/x= (3/9) -1/9 2/x=2/9 x=9

OpenStudy (anonymous):

Thank you so much @bere1 :)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

\[\frac{2}{x}+\frac{1}{9}=\frac{1}{3}\] multiplying and dividing by 9 and x resp. to terms of LHS \[\frac{2*9}{x*9}+\frac{1*x}{9*x}=\frac{1}{3}\] \[\frac{18}{9x}+\frac{x}{9x}=\frac{1}{3}\] \[\frac{18+x}{9x}=\frac{1}{3}\] cross-multiply \[3*(18+x)=9x\] \[54+3x=9x\] subtracting 3x from both sides \[54=6x\] dividing by 6 both sides \[9=x\]